OpenStudy (sro03):

Hi How can I change the Cartesian integral into an polar integral? Thanks in advance!

OpenStudy (sro03):

I have an exampel her: \[\int\limits_{0}^{2}\int\limits_{0}^{\sqrt{2x-x^2}}\sqrt{x^2+y^2} dy dx\]

hartnn (hartnn):

In general, plug in \(x = r\cos \theta \\ y = r\sin \theta\)

hartnn (hartnn):

here, \(r^2= x^2+y^2\)

hartnn (hartnn):

find what upper limit 2x-x^2 equals

hartnn (hartnn):

\(dy dx = r dr d\theta \)

OpenStudy (sro03):

But how can I know that the upper limit 2x-x^2 is in polar?

hartnn (hartnn):

y = 0 to y = \(\sqrt{2x-x^2}\) << this is a circle. can you plot it?

OpenStudy (sro03):

you mean that i should convert the upper limit to polar, right?

OpenStudy (sro03):

Yes I just have plot it now, and it's a half circel

hartnn (hartnn):

yes, \(0 \le x \le 2 \\ 0\le y \le \sqrt{2x-x^2}\) is equivalent to \( 0 \le \theta \le \pi/2 \\ 0\le r\le 2 \) makes sense?

hartnn (hartnn):

wait, my limits for 'r' aren't accurate

hartnn (hartnn):

since we have the circle, \(x^2+y^2 -2x= 0 \\ (x-1)^2 +y^2 = 1^2\) we can take \(x-1 = r \cos \theta \\ y = r\sin \theta\)

hartnn (hartnn):

\(0 \le x \le 2 \\ 0\le y \le \sqrt{2x-x^2} \) is now equivalent to \(0\le\ \theta \le \pi\\ 0\le r \le 1 \)

hartnn (hartnn):

you'll have to now find the value of \(\sqrt{x^2+y^2}\)

hartnn (hartnn):

\(\sqrt{(r\cos \theta +1)^2+(r\sin\theta)^2 } = \)

hartnn (hartnn):

that is r when we take x = r cos and y = r sin ... this is not what we have taken here

hartnn (hartnn):

\(\int_{0}^{\pi} \int_0^1 \sqrt{r^2+2r\cos\theta+1}r dr d\theta \)

hartnn (hartnn):

the conversion is done, do you also have to evaluate it?

hartnn (hartnn):

because evaluating that would be a challenge in polar form

OpenStudy (sro03):

Yes it shoudl also be evaluated :/ I use mathcad an maple.. but it can solve it :/

hartnn (hartnn):

ok, lets try \(u = r^2+ 2r \cos \theta +1 \\ du = (2r +2\cos \theta) dr \)

OpenStudy (kainui):

There are two competing choices for which polar coordinates to use, since we have the function centered at the origin and the upper bound centered at x=1. It turns out the one centered at the origin is easier, and we'll see in a sec why that's true: \[x=r \cos \theta\]\[y=r \sin \theta\] We plug these things into our set of bounds from the integral, \(y=0\), \(y=\sqrt{2x-x^2}\), \(x=0\), \(x=2\). Notice that we can rearrange that upper bound on y to look like this with minor algebra: \[x^2+y^2 = 2x\] Anywho, once we plug in these polar coordinates this becomes \[r^2 = 2 r \cos \theta\] in other words, \[r=2 \cos \theta\] Now if we pullback to this we have: \[\int_0^2 \int_0^{\sqrt{2x-x^2}} \sqrt{x^2+y^2} dydx = \int_0^\pi \int_0^{2 \cos \theta} r *rdrd\theta \] \[\int_0^\pi \int_0^{2 \cos \theta} r *rdrd\theta = \int_0^\pi \frac{1}{3} r^3 |_0^{2 \cos \theta} d \theta=\frac{8}{3} \int_0^\pi \cos^3 \theta d \theta\]

hartnn (hartnn):

nice! but wouldn't the limit be o to pi/2 here?

hartnn (hartnn):

the \(\int_0^\pi \cos^3 \theta = 0 \)

hartnn (hartnn):

must be a typo... \(\Large \int_0^{\pi/2} \int_0^{2 \cos \theta} r *rdrd\theta \\ \Large = \int_0^{\pi/2} \dfrac{1}{3} r^3 |_0^{2 \cos \theta} d \theta=\dfrac{8}{3} \int_0^{\pi/2} \cos^3 \theta d \theta = \dfrac{16}{3}\)

OpenStudy (kainui):

yeah haha

OpenStudy (sro03):

Really thanks to both of you because you help me!! Sorry i don't say it but I have the result, It should give 16/9 = 1,78

hartnn (hartnn):

oh yes, the cos integral is 2/3 so the final answer would be 16/9 indeed

OpenStudy (kainui):

Just out of curiosity, I wonder if the substitution \(u=x-1\) is worth looking at. It would turn the integral into (probably, I'm sloppy): \[\int_{-1}^1 \int_0^{\sqrt{1-u^2}} \sqrt{(u+1)^2+y^2} dy du\]

OpenStudy (kainui):

Anywho, I guess I'm just trying to help convey what sorta things are useful to look at in double and triple integrals. It's usually about choosing some choice of coordinates that leverages the symmetry of the situation. Usually you can find this out by taking all of the limits, writing them out and/or drawing them and then looking for where things are important. In this case we had two different interesting points and it looked like polar was better than rectangular at one of them. But overall there were 4 pretty decent possibilities. (Well, 3 since I guess the original one sucked which is why this question was asked in the first place and we needed to change lol)

OpenStudy (sro03):

Sorry, but I don't get it.. What should I change to 2/3 ? And sorry becuase of my bad english!

hartnn (hartnn):

\(\cos 3x = 4 \cos^3 x - 3\cos x\) \(\int_0^{\pi/2} \cos^3 x dx= \int_0^{\pi/2} \dfrac{\cos 3x + 3\cos x}{4}dx = (1/4)[\dfrac{\sin 3x}{3}+3\sin x ]_0^{\pi/2} \\= (1/4)[-1/3+3-0] = 2/3\) ^^

OpenStudy (sro03):

Can you help me to write a recipe / a step by step how to do list to solve or change cratesian integral into polar integral? :/